Suppose that one early morning when the temperature is 10 degree Celcius, a driver of an automobile gets his gasoline tank which is made of steel filled with 75 litres of gasoline which is also at 10 degree celcius. during the day the temperature rises to 30 degree celcius. How much gasoline overflows?


1
Expert's answer
2021-04-23T07:26:59-0400
"\Delta V_s=V\beta\Delta T\\\Delta V_s=(75)(36\cdot10^{-6})(30-10)=0.054\ L"

"\Delta V_g=V\beta\Delta T\\\Delta V_g=(75)(28.5\cdot10^{-4})(30-10)=4.275\ L\\\Delta V=4.275-0.054=4.221\ L"